OpenStudy (anonymous):

Simplify. (x – 3)(x2 – 4x – 7) A. x3 – 3x2 + 12x + 21 B. x3 – 7x2 + 19x + 21 C. x3 – 7x2 + 5x + 21 D. x3 + x2 – 5x – 21

OpenStudy (anonymous):

(x−3)((x(2)−4x)−7) =(x+(−3))(x(2)+(−4x)+(−7)) =(x)(x(2))+(x)(−4x)+(x)(−7)+(−3)(x(2))+(−3)(−4x)+(−3)(−7) =2x2−4x2−7x−6x+12x+21 =−2x2−x+21

OpenStudy (anonymous):

Now let's go down further:

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

((−2)(x)(2)−x)+21 =(−2)(x)(2)+(−x)+21 =(−5x)+21 =−5x+21

OpenStudy (anonymous):

Might be C

OpenStudy (anonymous):

alright thanks!

OpenStudy (whpalmer4):

Not only does @Saphira_Shapiro do the work for you, she gets the answer wrong. Maybe you ought to learn how to do it yourself :-) the x2 represents x*x or x^2, not x(2) or x*2 the x3 represents x*x*x or x^3, not x(3) or x*3

OpenStudy (anonymous):

@whpalmer4 If you have a problem, keep it to yourself. I learn how to solve that way. I am sorry you don't.

OpenStudy (anonymous):

I may not be Einstein but sometimes I get them correct too.

OpenStudy (whpalmer4):

Again, see the Code of Conduct: "Give Help, Not Answers"

OpenStudy (anonymous):

I have the same problem can someone help me understand it? and answer it?